a) Price of the Tablet (P610 Samsung Galaxy Tab-S6 Lite Octa-Core 4GB Ram 64GB Storage 

10.1 Inch Wifi with S Pen) increases from PKR 57000 to PKR 66000. Due to this change in 

price

1. What will happen with the Quantity Demanded?

2. Draw a demand curve accordingly.

3. What will happen with the Quantity Supplied?

4. Draw a supply curve accordingly?

b) In response to the above price change if quantity demanded changes 40% and quantity 

supplied changes 20%

1. Find the elasticity of demand and explain this change.

2. Find the elasticity of supply and explain this change.


1
Expert's answer
2021-04-26T19:30:10-0400

1) quantity demanded will decrease

2)




3) it will increase

4)





b) "\Delta Q_d = 40%"

"E_d = \frac{\Delta Q_d}{\frac{66000-57000}{57000}} = \frac{0,4}{\frac{9}{57}} = 2,53"

"E_s = \frac{0,2}{\frac{9}{57}} = 1,26"

The response to the price change from consumers is higher then from suppliers which means that the demand curve is more elastic than the supply curve

You

12:26


1) quantity demanded will decrease

2)




3) it will increase

4)





b) "\Delta Q_d = 40%"

"E_d = \frac{\Delta Q_d}{\frac{66000-57000}{57000}} = \frac{0,4}{\frac{9}{57}} = 2,53"

"E_s = \frac{0,2}{\frac{9}{57}} = 1,26"

The response to the price change from consumers is higher then from suppliers which means that the demand curve is more elastic than the supply curve

You

12:26


1) quantity demanded will decrease

2)




3) it will increase

4)





b) "\Delta Q_d = 40%"

"E_d = \frac{\Delta Q_d}{\frac{66000-57000}{57000}} = \frac{0,4}{\frac{9}{57}} = 2,53"

"E_s = \frac{0,2}{\frac{9}{57}} = 1,26"

The response to the price change from consumers is higher then from suppliers which means that the demand curve is more elastic than the supply curve